# A hyperbolic $4$-orbifold with underlying space $\mathbb{P}^{2}$

### v1

@@ 1 Introduction
The proof, given in Section 2, is giving by producing a triangulation of $\mathbb{P}^{2}$ that can be realized by $60$ copies of a particular compact Coxeter simplex in $\mathbb{H}^{4}$. The orbifold locus is described in Section 3. I also found examples with underlying space the product $\mathbb{S}^{2}\times\mathbb{S}^{2}$ of two spheres by similar methods, and Bruno Martelli described a convincing heuristic to me that one should also be able to construct a hyperbolic $4$-orbifold with underlying space the $4$-torus. The reason for focusing on $\mathbb{P}^{2}$ for this note is the following consequence, which is perhaps surprising in comparison with the fact that closed hyperbolic $4$-manifolds have signature zero by the Hirzebruch signature theorem [6, §19].
@@ Acknowledgments
I thank for Bruno Martelli for conversations related to this paper and his observation that this example was particularly interesting because the signature of the underlying space is nonzero. This material is based upon work supported by Grant Number DMS-2203555 from the National Science Foundation.
@@ 2 A simplicial decomposition of $\mathbb{P}^{2}$
The simplicial complex described in Tables 1 and 2 is a manifold homeomorphic to $\mathbb{P}^{2}$.
@@
The computational software Regina [1] can take a triangulation described as in this section and compute the homology groups and fundamental group. Regina quickly verifies that the given triangulation gives a simply connected manifold with $H_{2}(M)\cong\mathbb{Z}$, which must be $\mathbb{P}^{2}$ by Freedman’s theorem [2]. In fact, several iterations of Regina’s simplification algorithm quickly simplify the given triangulation to the four-simplex triangulation of $\mathbb{P}^{2}$ contained in Regina’s database of example manifolds. ∎
@@
By the Poincaré polyhedron theorem [7, Thm. 13.5.3], it suffices to show that the number of simplices around of type $(a_{0}\ a_{1}\ a_{2})$ divides $2\pi/\theta(a_{0}\ a_{1}\ a_{2})$, where $\theta(a_{0}\ a_{1}\ a_{2})$ is the dihedral angle of $\mathcal{S}$ at that triangle. Specifically,
[…]
and all other dihedral angles are $\pi/2$. An accounting of the the triangles with label (0 1 2)012(0\ 1\ 2)( 0 1 2 ) is in Table 3, and every degree is divisible by $8$, as required; interior points of a triangle with degree $8$ have neighborhoods isometric to $\mathbb{H}^{4}$, and if the degree is $d<8$, then any interior point is contained in the orbifold locus with weight $8/d$. Every triangle in the simplicial decomposition with the same labeling as a triangle of $\mathcal{S}$ with dihedral angle $\pi/3$ has exactly six simplices that meet at the given triangle, which makes the identification space locally isometric to $\mathbb{H}^{4}$ around points in the interior of those triangles. For the remaining triangles, where $\mathcal{S}$ has dihedral angle $\pi/2$, there are exactly four simplices around each triangle except those enumerated in Table 4, which have only two; those with four simplices have neighborhoods isometric to $\mathbb{H}^{4}$ and those with two are part of the orbifold locus and have weight two. This completes all the checks necessary to apply the Poincaré polyhedron theorem. ∎
@@
$2$ $e_{33},e_{50},e_{38}$ $v_{7},v_{6},v_{8}$
Table 3: Triangles of type (0 1 2)012(0\ 1\ 2)( 0 1 2 )
@@
(1 2 4)124(1\ 2\ 4)( 1 2 4 ) $e_{45},e_{47},e_{49}$ $v_{9},v_{2},v_{4}$
Table 4: Triangles with dihedral angle $\pi/2$ and degree $2$
@@ 3 The orbifold locus
$v_{1}$ $\Delta_{0}(1)$ $\Delta_{1}(1)$ $\Delta_{2}(1)$ $\Delta_{4}(1)$ $\Delta_{6}(1)$ $\Delta_{10}(1)$ $6$
@@
$v_{9}$ $\Delta_{31}(1)$ $\Delta_{42}(1)$ $\Delta_{43}(1)$ $\Delta_{52}(1)$ $\Delta_{53}(1)$ $\Delta_{58}(1)$ $6$
@@
Figure 1: The (0 1 2)012(0\ 1\ 2)( 0 1 2 ) orbifold locus
[…]

### v2

@@ 1 Introduction
The proof, given in Section 2, is giving by producing a simplicial decomposition of $\mathbb{P}^{2}$ that can be realized by $60$ copies of a particular compact Coxeter simplex in $\mathbb{H}^{4}$. The orbifold locus is described in Section 3. I also found examples with underlying space the product $\mathbb{S}^{2}\times\mathbb{S}^{2}$ of two spheres by similar methods, and Bruno Martelli described a convincing heuristic to me that one should also be able to construct a hyperbolic $4$-orbifold with underlying space the $4$-torus. The reason for focusing on $\mathbb{P}^{2}$ for this note is the following consequence, which is perhaps surprising in comparison with the fact that closed hyperbolic $4$-manifolds have signature zero by the Hirzebruch signature theorem [7, §19].
@@ Acknowledgments
I thank for Bruno Martelli for conversations related to this paper and his observation that this example was particularly interesting because the signature of the underlying space is nonzero. I also thank Jonathan Spreer for comments on smooth structures. The author acknowledges support from the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and LabEx CARMIN (ANR-10-LABx-59-01). This material is based upon work supported by NSF grants DMS-2203555 and DMS-2506896, along with award SFI-MPS-TSM-00014184 from the Simons Foundation. Part of this research was performed while the author was visiting the Mathematical Sciences Research Institute (MSRI), now becoming the Simons Laufer Mathematical Sciences Institute (SLMath), which is supported by the National Science Foundation (Grant No. DMS-1928930).
@@ 2 A simplicial decomposition of $\mathbb{P}^{2}$
The simplicial complex described in Tables 1 and 2 is a PL manifold that is smoothable to the standard smooth structure on $\mathbb{P}^{2}$.
@@
The computational software Regina [2] can take a simplicial decomposition described as in this section and compute the homology groups and fundamental group. Regina quickly verifies that the given simplicial decomposition gives a simply connected manifold with $H_{2}(M)\cong\mathbb{Z}$, which must be $\mathbb{P}^{2}$ by Freedman’s theorem [3]. In fact, several iterations of Regina’s simplification algorithm quickly simplify the given simplicial decomposition to the four-simplex simplicial decomposition of $\mathbb{P}^{2}$ contained in Regina’s database of example manifolds. As described by Burke, Burton, and Spreer [1, §4.4], this simplicial decomposition can be obtained by applying Regina’s standard simplifications, which do not change the underlying smooth structure, from a standard Kirby diagram for $\mathbb{P}^{2}$ associated with the standard smooth structure. This proves the theorem. ∎
@@
By the Poincaré polyhedron theorem [8, Thm. 13.5.3], it suffices to show that the number of simplices around a vertex of type $(a_{0}\ a_{1}\ a_{2})$ divides $2\pi/\theta(a_{0}\ a_{1}\ a_{2})$, where $\theta(a_{0}\ a_{1}\ a_{2})$ is the dihedral angle of $\mathcal{S}$ at that triangle. Specifically,
$\theta(0\ 1\ 2)$ $=\frac{\pi}{4}$ $\theta(0\ 1\ 4)$ $=\frac{\pi}{3}$ $\theta(0\ 3\ 4)$ $=\frac{\pi}{3}$
$\theta(2\ 3\ 4)$ $=\frac{\pi}{3}$ $\theta(1\ 2\ 3)$ $=\frac{\pi}{3}$
and all other dihedral angles are $\pi/2$. An accounting of the the triangles with label (0 1 2)(0\ 1\ 2) is in Table 3, and every degree is divisible by $8$, as required; interior points of a triangle with degree $8$ have neighborhoods isometric to $\mathbb{H}^{4}$, and if the degree is $d<8$, then any interior point is contained in the orbifold locus with weight $8/d$. Every triangle in the simplicial decomposition with the same labeling as a triangle of $\mathcal{S}$ with dihedral angle $\pi/3$ has exactly six simplices that meet at the given triangle, which makes the identification space locally isometric to $\mathbb{H}^{4}$ around points in the interior of those triangles. For the remaining triangles, where $\mathcal{S}$ has dihedral angle $\pi/2$, there are exactly four simplices around each triangle except those enumerated in Table 4, which have only two; those with four simplices have neighborhoods isometric to $\mathbb{H}^{4}$ and those with two are part of the orbifold locus and have weight two. This completes all the checks necessary to apply the Poincaré polyhedron theorem. ∎
@@ 22 $e_{33},e_{50},e_{38}$ $v_{7},v_{6},v_{8}$
22 $e_{33},e_{50},e_{38}$ $v_{7},v_{6},v_{8}$
Table 3: Triangles of type (0 1 2)(0\ 1\ 2)
@@
(1 2 4)(1\ 2\ 4) $e_{45},e_{47},e_{49}$ $v_{9},v_{2},v_{4}$
Table 4: Triangles with dihedral angle $\pi/2$ and degree 22
@@ 3 The orbifold locus
$v_{1}$ $\Delta_{0}(1)$ $\Delta_{1}(1)$ $\Delta_{2}(1)$ $\Delta_{4}(1)$ $\Delta_{6}(1)$ $\Delta_{10}(1)$ 66
@@
$v_{9}$ $\Delta_{31}(1)$ $\Delta_{42}(1)$ $\Delta_{43}(1)$ $\Delta_{52}(1)$ $\Delta_{53}(1)$ $\Delta_{58}(1)$ 66
@@
Figure 1: The (0 1 2)(0\ 1\ 2) orbifold locus
$v_{0}$$v_{2}$$v_{9}$$v_{1}$$v_{3}$$v_{4}$$v_{7}$$v_{8}$$t_{86}$$t_{129}$$t_{76}$$t_{126}$$e_{0}$$e_{4}$$e_{45}$$e_{44}$$t_{28}$$t_{1}$$t_{23}$$t_{132}$$t_{97}$$t_{131}$$C$ (orbifold weight 44)